OpenStudy (anonymous):

Hi, can someone please help me obtain the second derivative of: y=x^2 + x^(1/2)? I get the first as 2X + (1/2)X^(-1/2), but I am stumped on the 2nd. Thanks.

OpenStudy (anonymous):

The first is indeed right, the second is this: 2+(1/2)*(-1/2)x^(-3/2)=2-(1/4)x^(-3/2)

OpenStudy (anonymous):

\[f(x)=x^2+\sqrt{x}\] \[f'(x)=2x+\frac{1}{2\sqrt{x}}\] \[f''(x)=2 -\frac{1}{4\sqrt{x^3}}\]

OpenStudy (anonymous):

what thomas said. keep using power rule, in particular for \[\frac{1}{2\sqrt{x}}=\frac{1}{2}x^{-\frac{1}{2}}\]

OpenStudy (anonymous):

so \[\frac{d}{dx}\frac{1}{2}x^{-\frac{1}{2}}=-\frac{1}{2}\times \frac{1}{2}x^{-\frac{1}{2}-1}\]

OpenStudy (anonymous):

\[=-\frac{1}{4}x^{-\frac{3}{2}}=-\frac{1}{4\sqrt{x^3}}\]\]

OpenStudy (anonymous):

thanks for the help. this was a multiple choice question given by my Prof. for extra credit and the choices are: a)[(2X ^{3/2}-1)/ x^{3/2}] b)[(8x^{3/2}+1)/4x^{3/2}] c)[(8x ^{3/2}-1)/4x ^{3/2}] d)[(2x ^{3/2}+1)/x ^{3/2}] Thanks again.